# thanks: Dedicated to Prof. Jafar Sadeghi in honor of his 60th jubilee

### v1

@@ Abstract
In this paper, we study the universe in the context of $f(T)$ gravity with flat-FRW background and consider it to be a combination of three components: baryonic matter, dark matter, and dark energy. We consider the new agegraphic as an alternative for dark energy and the Bose-Einstein condensation (BEC) as an alternative for dark matter. After that, first we obtain the Friedman equations and then we obtain the continuity equations in the presence of the interaction term between the dark components of the universe, so that the interaction term is considered as the flow of energy from dark matter to dark energy. In what follows, we plot the variation of the cosmological parameters of dark energy in terms of the redshift parameter by using of the power-law cosmology. Finally, we investigate the evolution and stability of the universe and report the values of the density parameters of the universe components that confirm the present model with observational data.
Equation of state parameter; Dark energy; Dark matter; Bose-Einstein condensate; $f(T)$ gravity.
@@ I Introduction
One of these models is the teleparallel gravity model, which was first proposed by Einstein to unify electromagnetism and gravity. In this theory, space-time is characterized by a linear connection without curvature and with twist by a metric tensor field called the dynamical tetrad field. Therefore, the geometry of general relativity, which is based on the Levi-Civita connection for a torsion-free curvature connection, gives way to the geometry of the parallel trap theory, which is based on the linear Weitzenböck connection in a curvature-free torsion connection. [48, 49, 50, 51, 53, 52, 54, 55].
@@
On the other hand, another dark part of the universe is dark matter, which is not visible in any spectrum of electromagnetic radiation, it is an important topic in cosmology. Gravitational lensing and the cosmic microwave background are observational evidence of dark matter, which are calculated from the difference between the mass calculated for the giant celestial bodies [56, 57, 58, 59]. Therefore, this experimental evidence shows that dark matter causes the gravitational effect between the formation and evolution of galaxies and clusters and large-scale structures. For this purpose, we use the Bose-Einstein condensation model to understand the concept of dark matter. Bose-Einstein condensation occurs in boson-diluted gases, where a large number of bosons simultaneously occupy a ground quantum state in a system. A number of papers have addressed the connection between BEC and dark matter. Therefore, we follow the BEC model as one of the models that can investigate dark matter [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 77, 79, 80, 81, 82].
Therefore, in this work, we want to examine the evolution of the universe in the context of $f(T)$ gravity with equivalence BEC and the new agegraphic, respectively, with dark matter and dark energy. This study became the motivation to reach a new approach to understanding the universe by reconstructing this model in terms of the redshift parameter and replacing the dark parts with BEC and the new agegraphic.
@@
In Sec. II, we imply the fundamental formation of $f(T)$ gravity in the flat-FRW metric. In Sec. III, we consider an interacting model between the dark parts of the universe. In Sec. IV, we review the model of the new agegraphic. In Sec. V, we consider BEC as dark matter and obtain a equation of state (EoS) for dark matter. In Sec. VI, we reconstruct the model in terms of the redshift parameter, and correspond the dark energy part with the new agegraphic. Finally, in Sec. VII, we will present a brief explanation for the present model.
II $f(T)$ gravity background
Here, we intend to review modified teleparallel gravity by $f(T)$ gravity background. Therefore, instead of the Levi-Civita connection in general relativity, we consider the Weitzenböck connection in teleparallel gravity. For this purpose, the corresponding geometry for each point $x^{\mu}$ in a manifold is defined based on the tetrad field $e_{i}(x^{\mu})$ as the basis vectors at same point, which will have $e_{i}.e_{j}=\eta_{ij}$, where the sign of Minkowski spacetime is $\eta_{ij}=diag(+1,-1,-1,-1)$. However, we can convert metric tensor $g_{\mu\nu}$ to the tetrad field $e^{i}_{\,\,\mu}$ as $g_{\mu\nu}=\eta_{\mu\nu}e^{i}_{\,\,\mu}e^{i}_{\,\,\nu}$, which the relations $e_{i}\,^{\mu}e^{i}\,_{\nu}=\delta_{\nu}^{\mu}$ and $e_{i}\,^{\mu}e^{j}\,_{\mu}=\delta_{i}^{j}$ are used. Note that Greek and Latin letters represent space-time components and tangent space-time components, respectively.
Nevertheless, the Weitzenböck connection is written down as
$\Gamma^{\lambda}_{~{}\mu\nu}=e_{i}^{~{}\lambda}\partial_{\mu}e^{i}_{~{}\nu}=-e^{i}_{~{}\mu}\partial_{\nu}e_{i}^{~{}\lambda},$ (1)
based on this, the torsion tensor ${T^{\lambda}}_{\mu\nu}$, the torsion scalar $T$, the antisymmetric tensor $S_{\lambda}^{\mu\nu}$, and the contortion tensor $K_{\lambda}^{\mu\nu}$ are
${T^{\lambda}}_{\mu\nu}=\Gamma^{\lambda}_{~{}\mu\nu}-\Gamma^{\lambda}_{~{}\nu\mu}={e_{a}}^{\lambda}\,\left({{\partial_{\mu}}{e^{a}}_{\nu}-{\partial_{\nu}}{e^{a}}_{\mu}}\right),$ (2a)
$T={S_{\lambda}}^{\mu\nu}\,{T^{\lambda}}_{\mu\nu},$ (2b)
$S_{\lambda}^{\mu\nu}=\frac{1}{2}\,\left({K_{\lambda}^{\mu\nu}+\delta_{\lambda}^{\mu}\,T_{\alpha}^{\alpha\nu}-\delta_{\lambda}^{\nu}\,T_{\alpha}^{\alpha\mu}}\right),$ (2c)
$K_{\lambda}^{\mu\nu}=-\frac{1}{2}\,\left({{T^{\mu\nu}}_{\lambda}-T_{\lambda}^{\nu\mu}-T_{\lambda}^{\mu\nu}}\right),$ (2d)
Now in this paper, we intend to explore $f(T)$ gravity model by the following action
@@ II $f(T)$ gravity background
where are $\kappa^{2}=8\pi G$ and $e=det\left(e^{i}_{\,\,\mu}\right)$, also $\mathcal{L}_{m}$ is the matter Lagrangian density. As we know in physics, the principle of least action is the variation principle of a system, so that by varying the above action with respect to a the tetrad field, we can obtain the field equation as
$e^{-1}\partial_{\mu}\left(e\,e^{\lambda}_{a}\,S_{\lambda}^{~{}\mu\nu}\right)f_{T}-e_{a}^{\rho}\,T^{\lambda}_{~{}\mu\rho}S_{\lambda}{}^{\nu\mu}f_{T}+e^{\lambda}_{a}S_{\lambda}{}^{\mu\nu}\,\partial_{\mu}(T)f_{TT}+\frac{1}{4}e_{a}^{\nu}f=\frac{1}{2}\kappa^{2}e_{a}^{\lambda}\mathcal{T}_{\lambda}{}^{\nu},$ (4)
where the indices $T$ and $TT$ indicate the first and second order derivatives with respect to torsion scalar, and $\mathcal{T}_{\lambda}{}^{\nu}$ is the energy-momentum tensor. Therefore, the Ricci tensor $R_{\mu\nu}$ and the Ricci scalar $R$ obtain in the following form
$R_{\mu\nu}=-\nabla^{\rho}S_{\nu\rho\mu}-g_{\mu\nu}\nabla^{\rho}T^{\sigma}_{~{}\rho\sigma}-S^{\rho\sigma}_{~{}~{}\mu}K_{\sigma\rho\nu},$ (5a)
$R=-T-2\nabla^{\mu}\left(T^{\nu}_{~{}\mu\nu}\right).$ (5b)
@@
III Interacting model in $f(T)$ gravity background
@@
where $a(t)$ is the scale factor. The tetrad field is obtained as $e_{\mu}^{a}=\mathrm{diag}(1,a,a,a)$ or $e^{\mu}_{a}=\mathrm{diag}(1,a^{-1},a^{-1},a^{-1})$, and then the torsion scalar becomes:
@@
where indices $bm$, $dm$ and $de$ are introduced as the contents of baryonic matter, dark matter and dark energy. Hence, we can obtain the total continuity equation from Eqs. (10) as
@@
where $Q$ is the interaction term. The interaction term appears in the universe when there is a flow of energy between the components of the universe, so that the direction of the energy flow is from dark energy to dark matter and there is no interaction between baryonic matter and other components. With this approach, the exclusive interaction between dark matter and dark energy without the involvement of visible matter leads to fascinating cosmological consequences. Note that the interaction term is confirmed from the point of view of the second law of thermodynamics and its value is positive [83]. From the continuity equation, we find that the unit of term $Q$ is equal to the product of the Hubble parameter in the energy density. Therefore, in this work, we take the corresponding interaction term as $Q=3b^{2}H\rho_{dm}$ in which $b$ is intensity of the energy flow transfer between the corresponding components of the universe. From Eqs. (13a) and (13b) we find
@@
where $\omega_{bm}=p_{bm}/\rho_{bm}$ is Equation of State (EoS) of baryonic matter which is a constant, and $\rho_{bm_{0}}$ is the present dark energy of baryonic matter.
@@
where $\omega_{m}=p_{m}/\rho_{m}$ is as the matter EoS parameter which its value is a constant. Also, we can write down the EoS of dark energy as follows:
$\omega_{de}=\frac{p_{de}}{\rho_{de}}=-1-\frac{2\dot{H}\partial_{T}f-24H^{2}\dot{H}\partial_{TT}f+\kappa^{2}\rho_{m}\,(1+\omega_{m})}{6H^{2}\partial_{T}f+\frac{1}{2}f-\kappa^{2}\rho_{m}},$ (16)
where the obtained EoS of dark energy is a modified form, this means that if we choose $f(T)=T$ without matter, it becomes $\omega=-1-\frac{2\dot{H}}{3H^{2}}$, which is same as the standard form in general relativity. Then, the obtained EoS shows us that the nature of the universe depends on $f(T)$ gravity and matter.
IV The new agegraphic dark energy model
In this section, we give a general overview of the subject of the agegraphic. As we know the basis of quantum mechanics is based on the fact that things that can be measured have a special meaning in physics. Therefore, experimental observations can be the result of thoughts related to the specific science. In that case, the measurement of physical quantities in general relativity and quantum mechanics is different from each other in the sense that in the former, quantities are measured without any restrictions, but in the latter, the measurement is done with restrictions and discretely. Nevertheless, the theory of gravity that incorporates the principles of quantum mechanics is known as the quantum theory of gravity. Although the theory of quantum gravity is not fully established, some concepts based on the principles of quantum gravity are presented. This means that when gravitational fields such as the inflation period and the nature of black holes and quantum effects in gravity are very strong, they can be investigated using quantum gravity theory. Therefore, the phenomenology of space-time quantum fluctuations, which is explored by the theory of quantum gravity, is introduced as an agegraphic model. Accordingly, using the concept of space-time quantum fluctuations, the distance $t$ (with the speed of light $c=1$) in Minkowski space-time was determined according to the uncertainty principle with the following accuracy
$Eq16$ (17)
δ t=λ t_p^2/3t^1/3, where $t_{{p}}$ is Planck’s reduced time and $\lambda$ is a dimensionless constant of order unity [84, 85, 86, 87, 88]. The energy density obtained from the agegraphic model is equal to $\rho_{{DE}}\approx{\frac{1}{{\kappa}^{2}{t}^{2}}}$, where $t$ is the age of the universe. By substituting the age of the universe $t$ in the energy density of the agegraphic dark energy model with conformal time $\tau$, the energy density of the new agegraphic dark energy model is obtained as follows:

### v2

@@ Abstract
In this paper, we investigate the cosmic evolution within the framework of $f(T)$ gravity using a flat-FRW background and model the universe as consisting of three components: baryonic matter, dark matter, and dark energy. We consider the new agegraphic (NA) as an alternative for dark energy and the Bose-Einstein condensation (BEC) as an alternative for dark matter. After that, first we obtain the Friedman equations and then we obtain the continuity equations in the presence of the interaction term between the dark components of the universe, where the interaction term represents the energy flow from dark matter to dark energy. In what follows, we plot the variation of the cosmological parameters of dark energy in terms of the redshift parameter by using the power-law cosmology. Finally, we investigate the evolution and stability of the universe and report the values of the density parameters of the universe components which confirm the present model with observational data.
Equation of state parameter; New agegraphic dark energy; Bose-Einstein condensate dark matter; $f(T)$ gravity.
@@ I Introduction
In recent years, modified theories of gravity have received widespread attention as promising frameworks for explaining the acceleration of the universe in recent times, without the need for dark energy of unknown nature. One of these models is the teleparallel gravity model, which was first proposed by Einstein to unify electromagnetism and gravity. In this theory, space-time is characterized by a linear connection without curvature and with twist by a metric tensor field called the dynamical tetrad field. Therefore, the geometry of general relativity, which is based on the Levi-Civita connection for a torsion-free curvature connection, gives way to the geometry of the teleparallel theory, which is based on the linear Weitzenböck connection in a curvature-free torsion connection. [48, 49, 50, 51, 53, 52, 54, 55]. Among these theories, $f(T)$ gravity, a generalization of teleparallel gravity in which the Lagrangian is defined as an arbitrary function of the torsion scalar $T$, offers a geometric structure and second-order field equations that distinguish it from higher-order theories such as $f(R)$. This torsion-based framework enables new cosmic dynamics and has been used in inflationary models, dark energy scenarios, and cosmic acceleration [56, 57].
@@
On the other hand, the nature of dark matter remains one of the great mysteries of cosmology. While the standard cold dark matter (CDM) model has performed well on large scales, it faces challenges on galactic and subgalactic scales, such as the core-tip problem and the lack of satellites. An attractive alternative is the Bose-Einstein condensate (BEC) dark matter model, which models dark matter as a quantum fluid of ultralight bosons that condense on cosmic scales. This approach naturally introduces quantum pressure and internal potential into the model and can help smooth halo profiles and resolve small anomalies. However, dark matter is not visible in any spectrum of electromagnetic radiation, but it is an important topic in cosmology. Gravitational lensing and the cosmic microwave background are observational evidence of dark matter, as inferred from discrepancies between visible mass and gravitational effects observed in galaxy clusters. [58, 59, 60, 61]. Therefore, this experimental evidence shows that dark matter causes the gravitational effect between the formation and evolution of galaxies and clusters and large-scale structures. For this purpose, we use the Bose-Einstein condensation model to understand the concept of dark matter. Bose-Einstein condensation occurs in boson-diluted gases, where a large number of bosons simultaneously occupy a ground quantum state in a system. A number of papers have addressed the connection between BEC and dark matter. Therefore, we consider the BEC model as a viable candidate for describing dark matter [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 79, 81, 82, 83, 84].
The NA dark energy model is based on the quantum uncertainty principle and the energy-time relation in quantum mechanics. In this model, the dark energy density is related to the inverse square of cosmic time, but unlike the original version, conformal time is used instead of cosmic time. This modification makes the model more consistent with observational data and the dynamical behavior of the universe. We will describe the NA in more detail in Section IV [85, 86, 87, 88, 89].
The combination of the $f(T)$ theory of gravity, Bose-Einstein condensed dark matter (BEC), and the NA dark energy model provides a new perspective for describing the universe. In the $f(T)$ theory, the curvature of space-time is replaced by torsion, which allows for geometric modifications to the field equations. On the other hand, BEC dark matter, with its quantum pressure and wave behavior, can shapes cosmic structures differently from classical cold dark matter. The NA model also describes dark energy using the time scale of the universe and the quantum uncertainty principle, which is naturally consistent with the properties of the BEC.
By combining these three components, it is possible to build a model in which the geometric torsion $f(T)$ affects the behavior of quantum dark matter, while dark energy appears dynamically via the agegraphic time scale. This synergy not only allows for a description of the cosmic acceleration without the need for a cosmological constant, but can also have observable effects on gravitational lensing data, large-scale structure, and fluctuations in the cosmic microwave background.
The main advantage of this approach is in creating a coherent, multi-layered cosmological framework that takes into account both the geometric aspects (torsion), the quantum properties of dark matter, and the dynamics of dark energy, and this could lead to a deeper understanding of the nature of the dark matter components of the Universe.
In an effort to comprehensively understand the nature of the accelerating universe, this research is based on three complementary frameworks: NA dark energy, BEC dark matter, and $f(T)$ gravity. It is hypothesized that the interaction between these three components can provide a coherent and multi-layered picture of the evolution of the universe, such that the geometric torsion in f(T)f(T)gravity affects the dynamical behavior of BEC-type dark matter, while dark energy emerges in a quantum dynamical form and on a homomorphic time scale in the NA model. This theoretical equivalence not only allows for the description of cosmic acceleration without the need for the cosmological constant $\Lambda$, but also can yield results consistent with observational data such as the Hubble parameter, the cosmic microwave background radiation, and large-scale structures. Accordingly, the aim of this research is to reconstruct and analyze the behavior of the universe in the framework of $f(T)$ by considering BEC as dark matter and NA as dark energy. For this purpose, the Friedmann and continuity equations are extracted for each component and the effect of the interaction term $Q$ on the energy exchange between the dark sectors is investigated. Also, by applying power law cosmology and reconstructing the function $f(T)$ in terms of the torsion scalar from observational data, thermodynamic stability and the transition from the matter-dominated to the acceleration-dominated phase are investigated. Finally, the agreement of the obtained values of the density parameters $\Omega_{bm_{0}}$, $\Omega_{dm_{0}}$ and $\Omega_{de_{0}}$ with the observed values is a confirmation of the efficiency of the proposed model in describing the current universe as a dynamic, geometric and quantum system.
@@
In Sec. II, we review the fundamental formulation of $f(T)$ gravity in the flat-FRW metric. In Sec. III, we consider an interacting model between the dark parts of the universe. In Sec. IV, we review the models of the NA dark energy and BEC dark matter, and we respectively obtain the energy density of the NA dark energy and Equation of State (EoS) for BEC dark matter. In Sec. V, we reconstruct the model in terms of the redshift parameter and establish a correspondence between the dark energy and dark matter components. Finally, in Sec. VI, we will present a brief explanation for the present model.
II Covariant $f(T)$ gravity background
In this section, we intend to review modified teleparallel gravity by $f(T)$ gravity background within the covariant formulation. This formulation is able to guarantee local Lorentz invariance and frame independence. If we consider the model of $f(T)$ gravity only with the traditional pure-tetrad approach without explicitly introducing a spin connection, it violates the local Lorentz symmetry. That is, the investigation of the theory depends on the specific choice of tetrad, which can lead to different results with different choices. To avoid this problem, we include both the tetrad $e^{a}_{\mu}$ and a non-trivial spin connection $\omega^{\lambda}_{a\nu}$ in the theory, in order to define a covariant derivative and construct the torsion tensor, which can lead to Lorentz invariance. Therefore, the resulting equations are consistent with the physical concept, and are free of artifacts related to the tetrad frame.
Now, we start from the theory action of $f(T)$ gravity as follows:
@@ II Covariant $f(T)$ gravity background
where $\kappa^{2}=8\pi G$ is the Einstein gravitational constant, $e=det\left(e^{i}_{\,\,\mu}\right)$ is the determinant of the tetrad field $e^{i}_{\,\,\mu}$, and $\mathcal{L}_{m}$ is the matter Lagrangian density. In addition, we can transform metric tensor $g_{\mu\nu}$ into the tetrad field $e^{i}_{\,\,\mu}$ in the form of $g_{\mu\nu}=\eta_{\mu\nu}e^{i}_{\,\,\mu}e^{i}_{\,\,\nu}$ ($\eta_{ij}=diag(+1,-1,-1,-1)$ is Minkowski space-time), where the relations $e_{i}\,^{\mu}e^{i}\,_{\nu}=\delta_{\nu}^{\mu}$ and $e_{i}\,^{\mu}e^{j}\,_{\mu}=\delta_{i}^{j}$ hold. Note that Greek and Latin letters represent space-time components and tangent space-time components, respectively. The torsion scalar $T$ is defined as
$T={S_{\lambda}}^{\mu\nu}\,{T^{\lambda}}_{\mu\nu},$ (2)
where the torsion tensor ${T^{\lambda}}_{\mu\nu}$, the antisymmetric tensor $S^{\mu\nu}{}_{\lambda}$, and the contortion tensor $K^{\mu\nu}{}_{\lambda}$ are as follows:
[…]
$K^{\mu\nu}{}_{\lambda}=-\frac{1}{2}\,\left({{T^{\mu\nu}}_{\lambda}-{T^{\nu\mu}}_{\lambda}-{T_{\lambda}}^{\mu\nu}}\right),$ (3c)
The variation of the action with respect to the tetrad yields the Einstein’s field equations:
$e^{-1}\partial_{\mu}\left(e\,e^{\lambda}_{a}\,S_{\lambda}^{~\mu\nu}\right)f_{T}-e_{a}^{\rho}\,T^{\lambda}_{~\mu\rho}S_{\lambda}{}^{\nu\mu}f_{T}+e^{\lambda}_{a}S_{\lambda}{}^{\mu\nu}\,\partial_{\mu}(T)f_{TT}+\omega^{\lambda}_{~a\nu}S_{\lambda}{}^{\mu\nu}f_{T}+\frac{1}{4}e_{a}^{\nu}f=\frac{1}{2}\kappa^{2}e_{a}^{\lambda}\mathcal{T}_{\lambda}{}^{\nu}$, (4)
where the indices $T$ and $TT$ indicate the first and second order derivatives with respect to torsion scalar, and $\mathcal{T}_{\lambda}{}^{\nu}$ is the energy-momentum tensor.
In the covariant formulation, the spin connection $\omega^{\lambda}_{a\nu}$ is chosen such that the total curvature vanishes, i.e., $R^{\lambda}_{~b\mu\nu}(\omega)=0$. This condition ensures that the geometry remains teleparallel (For more details, refer to Ref. [90]).
@@
III the modified Friedmann equations and the continuity equations
@@
where $a(t)$ is the scale factor. The diagonal tetrad field is obtained as $e_{\mu}^{a}=\mathrm{diag}(1,a,a,a)$ or $e^{\mu}_{a}=\mathrm{diag}(1,a^{-1},a^{-1},a^{-1})$ that is compatible with a vanishing spin connection, making it a ”good tetrad” in the covariant sense. The torsion scalar is as
@@
where indices $bm$, dmdm and $de$ are introduced as the contents of baryonic matter, dark matter and dark energy, respectively. Hence, we can obtain the total continuity equation from Eqs. (9) as
@@
where $Q$ is the interaction term. The interaction term appears in the universe when there is a flow of energy between the components of the universe, so that the direction of the energy flow is from dark energy to dark matter and there is no interaction between baryonic matter and other components. With this approach, the exclusive interaction between dark matter and dark energy without the involvement of visible matter leads to fascinating cosmological consequences. Note that the interaction term is confirmed from the point of view of the second law of thermodynamics and its value is positive [91]. From the continuity equation, we find that the unit of term $Q$ is equal to the product of the Hubble parameter in the energy density. Therefore, in this work, we take the corresponding interaction term as $Q=3b^{2}H\rho_{dm}$ in which $b$ is intensity of the energy flow transfer between the corresponding components of the universe. It can be motivated by field-theoretic considerations, such as coupling between scalar fields or effective interactions in modified gravity. The presence of $Q$ allows for dynamical adjustment of energy densities, potentially resolving the cosmic coincidence problem and influencing the effective EoS.
From Eqs. (12a) and (12b) we find
@@
where $\omega_{bm}=p_{bm}/\rho_{bm}$ is EoS of baryonic matter which is a constant, and $\rho_{bm_{0}}$ is the present dark energy of baryonic matter.
@@
where $\omega_{dm}=p_{dm}/\rho_{dm}$ is as the matter EoS parameter whose value is constant. Also, we can write down the EoS of dark energy as follows:
$\omega_{de}=\frac{p_{de}}{\rho_{de}}=-1-\frac{2\dot{H}\partial_{T}f-24H^{2}\dot{H}\partial_{TT}f+\kappa^{2}\rho_{bm}\,(1+\omega_{bm})+\kappa^{2}\rho_{dm}\,(1+\omega_{dm})}{6H^{2}\partial_{T}f+\frac{1}{2}f-\kappa^{2}\rho_{bm}-\kappa^{2}\rho_{dm}},$ (15)
where the obtained EoS of dark energy is a modified form, this means that if we choose $f(T)=T$ in the absence of matter, it becomes $\omega=\frac{p_{tot}}{\rho_{tot}}=-1-\frac{2\dot{H}}{3H^{2}}$, which is same as the standard form in general relativity. Then, the obtained EoS shows us that the nature of the universe depends on $f(T)$ gravity and matter.
IV NA dark energy and BEC dark matter models
In this section, we will examine the two main dark components in the present cosmological model: NA dark energy and BEC dark matter. Both components are rooted in quantum mechanical principles and play a complementary role in describing cosmic acceleration in the modified gravity framework.
First, we give a general overview of the subject of the NA dark energy. As we know the basis of quantum mechanics is based on the fact that things that can be measured have a special meaning in physics. Therefore, experimental observations can be the result of thoughts related to the specific science. In that case, the measurements of physical quantities in general relativity and quantum mechanics are different from each other in the sense that in the former, quantities are measured without any restrictions, but in the latter, the measurements are restricted and occur in a discrete manner due to quantum effects. Nevertheless, the theory of gravity that incorporates the principles of quantum mechanics is known as the quantum theory of gravity. Although the theory of quantum gravity is not fully established, some concepts based on the principles of quantum gravity are presented. This means that when gravitational fields such as the inflation period and the nature of black holes and quantum effects in gravity are very strong, they can be investigated using quantum gravity theory. Therefore, the phenomenology of space-time quantum fluctuations, which is explored by the theory of quantum gravity, is introduced as an agegraphic model. Accordingly, using the concept of space-time quantum fluctuations, the distance $t$ (with the speed of light $c=1$) in Minkowski space-time was determined according to the uncertainty principle with the following accuracy
${Eq16}\delta\,t=\lambda\,{t_{{p}}}^{2/3}{t^{1/3}},$ (16)
where $t_{{p}}$ is Planck’s reduced time and $\lambda$ is a dimensionless constant of order unity [85, 86, 87, 88, 89]. The energy density obtained from the NA model is equal to $\rho_{{DE}}\approx{\frac{1}{{\kappa}^{2}{t}^{2}}}$, where $t$ is the age of the universe. By substituting the age of the universe $t$ in the energy density of the NA dark energy model with conformal time $\tau$, the energy density of the NA dark energy model is obtained as follows: